OpenStudy (anonymous):

Solve y=2xy+x^2 for y.

mathslover (mathslover):

I am solving this question in terms of x : y - 2xy = x^2 y(1-2x) = x^2 \(y = \frac{x^2}{1-2x} \) Actually it was already solved but I have solved again without having y in RHS ...